OpenStudy (curry):

Solve using contour integrals. integral [0,pi] of (d(theta))/(A-Bcos(theta))

OpenStudy (curry):

\[\int\limits_{0}^{\pi}\frac{ 1 }{ (A-Bcos(\theta)) }\]

OpenStudy (curry):

If you have another method, then that is fine also.

OpenStudy (amistre64):

is this simpler to line integrals?

OpenStudy (amistre64):

\[\int_C~ f(x,y)~ds\] \[\int\int~ f(x,y)~\sqrt{(x')^2+(y')^2}~dy~dx\]

OpenStudy (amistre64):

well, the page im looking at uses dt instead of dydx a spose since its a parametric

OpenStudy (curry):

Wait, if it was parametric, wouldn't the problem get more cmplicated?

OpenStudy (amistre64):

lol, at the moment i dont know

OpenStudy (curry):

hmmm, i tried taking the antiderivative, but that method isn't pretty here...

OpenStudy (amistre64):

http://www1.mat.uniroma1.it/~isopi/didattica/vc/Trapper.pdf this page may be useful, trying to review it

OpenStudy (solomonzelman):

I wouldn't take an anti derivative. Or perhaps on my paper, not to put it in here. I wish I could help. When I actually learn some math, I'll be hopefully able to help by then.

OpenStudy (curry):

hmmm, let me try it! can we both do it, so we can check each other's answer?

OpenStudy (curry):

so first of, would I make the substitution costheta = ((1/2) (z + 1/2))

OpenStudy (freckles):

I think the sub tan(x/2)=theta might work

OpenStudy (curry):

Um, i can't simply it enough, any more help?

OpenStudy (freckles):

\[\tan(\frac{x}{2})=\theta \]  \[\sin(x)=2 \sin(\frac{x}{2})\cos(\frac{x}{2})=2\frac{\theta}{\sqrt{\theta ^2 +1 }}\frac{1}{\sqrt{\theta ^2+1}}=\frac{2 \theta }{ \theta ^2+1} \\ \cos(x)=\cos^2(\frac{x}{2})-\sin^2(\frac{x}{2})=\frac{1}{\theta ^2 +1 } -\frac{\theta ^2}{\theta ^2 +1 } =\frac{1- \theta ^2 }{\theta ^2 +1} \\ \frac{1}{2} \sec^2(\frac{x}{2})dx=d \theta \\ dx =2 \cos^2(\frac{x}{2}) d \theta \\ dx =2 \frac{1}{\theta ^2 +1 } d \theta \\ dx=\frac{2 }{ \theta ^2 +1 } d \theta \] \[\int\limits_{}^{}\frac{1}{A-B \cos(x)} d x =\int\limits_{}^{}\frac{1}{A-B \cdot \frac{1-\theta ^2}{1+\theta ^2}} \frac{2}{\theta ^2 +1} d \theta \]